OpenStudy (anonymous):

Compute the following indefinite integral (7x^4)*(E^X^5)*dx

OpenStudy (anonymous):

\[\int\limits 7x^4 e ^{x ^{5}}dx\] put x^5=t \[5x^4 dx=dt,x^4~dx=\frac{ 1 }{ 5 }dt\] complete it.

OpenStudy (anonymous):

I'm really confused...I understand \[x^5=t\] and \[dt=5x^4dx\] but I don't understand \[x^4dx=1/5dt\]. This part is confusing on my printed notes as well and its the only part I seem to see.

OpenStudy (anonymous):

Ok. I just figured that piece out. But I'm confused on how to put it all together.

OpenStudy (anonymous):

I think I got it. This is really a lot of substitution essentially and keeping track of alot of little things. \[7(1/5dt(e^t))\] 7/5*e^(x^5)+c Thanks